# Interpret and Control Dense Retrieval with Sparse Latent Features

### v1

@@ Abstract
Hao Kang Carnegie Mellon University Pittsburgh, PA 15213 haok@andrew.cmu.edu                        Tevin Wang Carnegie Mellon University Pittsburgh, PA 15213 tevinw@andrew.cmu.edu                        Chenyan Xiong Carnegie Mellon University Pittsburgh, PA 15213 cx@andrew.cmu.edu
@@ 2 Methodology
Table 1: Reconstruction evaluation of sparse latent features and the reconstructed embeddings learned by our $k$-sparse autoencoder. MSE measures the embedding differences between original and reconstructed embeddings.
@@
(a) MsMarco (Rec.)
(b) MsMarco (Spr.)
(c) BEIR (Rec.)
(d) BEIR (Spr.)
Figure 2: Retrieval performance of reconstructed (Rec.) embeddings and the sparse latent features (Spr.) before and after the contrastive loss KLD is applied, on MsMarco and Beir.
@@ 3 Experiments
For generalizability on diverse retrieval tasks, we additionally evalute the sparse autoencoder on datasets from the Beir benchmark, such as TrecCovid, NaturalQuestions, and DBPediaEntity (Kwiatkowski et al., 2019; Hasibi et al., 2017; Thakur et al., 2021).
Figure 3: Frequency distribution comparison between bag-of-words and sparse latent features in MsMarco.
Interpretability Study. To assess interpretability, we generate N2G explanations (Foote et al., 2023). For each feature, we create a training set of 512 samples by selecting the highest-activating documents. We then perform forward passes on prefix sequences to extract activation sequences, which are input into N2G to construct trie representations for each feature. Gpt-4o is used to interpret each trie’s semantic meaning. More details on N2G can be found in Appendix B.
@@
Table 2: Examples of sparse latent features.
@@
4 Evaluation
Table 3: Features for the binary perspective query “What is the primary focus of a university education?” and the top result after dense retrieval on the reconstructed embeddings. Feature activations wer amplified by 0.5. B/A displays the number of documents related to the feature before and after the amplification on $k=5$ retrieval.
@@ 4 Evaluation
4.1 Retrieval Performance
4.2 Interpretability Study
As illustrated in Figure 3, the learned sparse latent features also follow Zipf’s law, but its distribution is less head-heavy. This is interesting as top-ranking features in the bag-of-words model are often common stop words, but the sparse latent features may skip these stop words and capture fine-grained and conceptually meaningful categories. Some of the feature examples explained by N2G are presented in Table 2 and more can be found in Appendix B.
4.3 Controllability Study
Figure 4: Improvement with retrieval scores on manipulated documents and queries by amplifying the relevant sparse latent features across different amounts.
Table 3 presents one example of controlling the retrieval results by manipulating the reconstructed query embeddings via the latent space. It shows that amplifying the targeted feature dimension effectively biases the retrieval results towards the corresponding perspective, i.e., “job” (84340) or “learning” (179723). This indicates that the learned faithful latent space provides a new mechanism to control the retrieval behavior which leads to many potential applications, for example, in enhancing the safety with human intervention in dense retrieval systems. Additional examples can be found in Appendix C.
@@ Appendix A Training Procedures
Appendix B Interpretability Study
This section presents the interpretability analysis for the latent features extracted by the $k$-sparse autoencoder using N2G explanations.
N2G provides an automated approach to interpreting the behavior of individual neurons by converting their activations into graph-based representations. It identifies the most relevant tokens that strongly activate a neuron and focuses on them by pruning the surrounding, less relevant context. This process isolates the essential patterns that contribute to the neuron’s activation.
In our analysis, we utilize the N2G approach to interpret latent features of the model. Sampled features from different parts of the frequency distribution (i.e. head, torso, tail) are shown in Table 5 along with their N2G explanations. Activated features and their associated semantic concepts for a subset of queries from MsMarco dataset are displayed in Table 6.
Appendix C Controllability Study

### v2

@@ Abstract
Hao Kang Carnegie Mellon University Pittsburgh, PA 15213 haok@andrew.cmu.edu                        Tevin Wang Carnegie Mellon University Pittsburgh, PA 15213 tevinw@andrew.cmu.edu                        Chenyan Xiong Carnegie Mellon University Pittsburgh, PA 15213 cx@cs.cmu.edu
@@ 2 Methodology
Table 1: Reconstruction evaluation of sparse latent features and the reconstructed embeddings learned by our $k$-sparse autoencoder from the Bge model. MSE measures the embedding differences between original and reconstructed embeddings. Results for the alternative MiniCPM embedding model can be found in Appendix E.
@@ 3 Experiments
For generalizability on diverse retrieval tasks, we additionally evalute the sparse autoencoder on datasets from the Beir benchmark, such as TrecCovid, NaturalQuestions, and DBPediaEntity (Kwiatkowski et al., 2019; Hasibi et al., 2017; Thakur et al., 2021). Additionally, we investigate the impact of the base embedding by applying our approach to an alternative embedding model, MiniCPM  [^3] https://huggingface.co/openbmb/MiniCPM-Embedding (Hu et al., 2024).
Interpretability Study. To assess interpretability, we generate N2G explanations (Foote et al., 2023). N2G provides an automated approach to interpret the behavior of individual neurons by converting their activations into graph-based representations. It identifies the most relevant tokens that strongly activate a neuron and focuses on them by pruning the surrounding, less relevant context. This process isolates the essential patterns that contribute to the neuron’s activation.
For each feature, we create a training set of 512 samples by selecting the highest-activating documents. We then perform forward passes on prefix sequences to extract activation sequences, which are input into N2G to construct trie representations for each feature. Gpt-4o-mini is used to interpret each trie’s semantic meaning.
@@
4 Evaluation
4.1 Retrieval Performance
(a) MsMarco (Rec.)
(b) MsMarco (Spr.)
Figure 2: Retrieval performance of reconstructed (Rec.) embeddings and the sparse latent features (Spr.) before and after the contrastive loss KLD is applied on MsMarco using Bge as the embedding model. Results on Beir can be found in Appendix B.
4.2 Interpretability Study
(a) Unigram Bag-of-Words
(b) Sparse Latent Features
Figure 3: Frequency distribution comparison between bag-of-words and sparse latent features in MsMarco using Bge as the embedding model. The high-frequency region is characterized by a small number of words that occur with extreme regularity, whereas the low-frequency region consists of a large proportion of words that appear only a limited number of times throughout the dataset.
Table 2: Examples of sparse latent features using Bge as the embedding model explained by N2G from different parts of the frequency distribution.
@@ 4.2 Interpretability Study
Table 3: Top activated features using Bge as the embedding model from the document “A few people reported that they paid their attorney as little as \$50 per hour, and a few reported paying as much as \$400 to \$650 per hour. But the vast majority paid between \$150 and \$350 per hour, with \$250 being the most commonly reported fee. The survey asked respondents about a number of things, including: 1 how much their divorce attorney charged per hour. 2 how much their divorce cost. 3 the number of issues that they resolved out of court and in court. 4 whether their spouse contested the case. 5 how long the divorce took from start to finish.”
Description from N2G
1. cost, pricing, expenses, rates, income
2. time, duration, sleep, hours, minutes
3. government, law, agencies, constitution, enforcement
4. tennis, courts, wimbledon, justices, decisions
5. health, anxiety, symptoms, stress, concerns
As illustrated in Figure 3, the learned sparse latent features also follow Zipf’s law, but its distribution is less head-heavy. This is interesting as top-ranking features in the bag-of-words model are often common stop words, but the sparse latent features may skip these stop words and capture fine-grained and conceptually meaningful categories. Representative feature examples extracted by N2G from different segments of the distribution are provided in Table 2, while the top activated features for a sampled document in MsMarco dataset are detailed in Table 3. Additional examples can be found in Appendix C.
4.3 Controllability Study
(a) Manipulated Document
(b) Manipulated Query
Figure 4: Improvement in retrieval scores on manipulated documents and queries by amplifying relevant sparse latent features across varying amounts using Bge as the embedding model. The x-axis is in logarithmic scale for better visualizing the trends since each step gets incremented by a factor of 2.
Table 4: Features for the binary perspective query “What is the primary focus of a university education?” and the top result after dense retrieval on the reconstructed embeddings using Bge as the embedding model. Feature activations were amplified by 0.5. B/A displays the number of documents related to the feature before and after the amplification on $k=5$ retrieval.
@@ 4.3 Controllability Study
Table 4 presents one example of controlling the retrieval results by manipulating the reconstructed query embeddings via the latent space. It shows that amplifying the targeted feature dimension effectively biases the retrieval results towards the corresponding perspective, i.e., “job” (84340) or “learning” (179723). This indicates that the learned faithful latent space provides a new mechanism to control the retrieval behavior which leads to many potential applications, for example, in enhancing the safety with human intervention in dense retrieval systems. Additional examples can be found in Appendix D.
@@ Appendix A Training Procedures
Appendix B Ablation Study
This section presents the ablation study, comparing models trained with MSE alone against those incorporating contrastive loss on the Beir dataset. The comparison is illustrated with Figure 5.
Appendix C Interpretability Study
In our interpretability analysis, we utilize the N2G approach to interpret latent features extracted by the autoencoder. Sampled features from different parts of the frequency distribution (i.e. head, torso, tail) are shown in Table 6 along with their N2G explanations. Activated features and their associated semantic concepts for a subset of queries from MsMarco dataset are displayed in Table 7.
Appendix D Controllability Study
@@
Appendix E Role of Base Embedding
This section explores the transferability of our method across different embedding models. As illustrated in Table 8, our approach demonstrates consistent performance when applied to the MiniCPM embedding. However, we observe a noticeable decline in retrieval accuracy when using the sparse autoencoder with $K=32$ active features. This reduction may be attributed to the significantly larger embedding dimension involved, which is three times the size of BgeBase. This increased dimensionality likely necessitates a greater number of active features to support the retrieval task. Additionally, the results of our interpretability analysis and controllability study, conducted using the MiniCPM embedding, are presented in Tables 9, 10, and 11.